OpenStudy (anonymous):

What is the answer to this proportion 4/12=y/9

Directrix (directrix):

Use the Means-Extremes Property (Cross Multiply) 4/12=y/9 12*y = 4*9 Solve for y @yalo

OpenStudy (dan815):

\[ 4+ 4+ 4 = 12\] 3 parts of 4 added up give you 12 so, Therefor just a 4 is 1/3rd of 12 \[4/12 = 1/3 = y/9\] What should y be to be 1/3rd of 9 Lets see can we break 9 into 3 parts and figure it out \[ 3+ 3+ 3 = 9\] so for y to be 1/3rd of 9 what should it be?